Eric invested his savings in two investment funds. The amount he invested in Fund A was 7000 less than the amount he invested in Fund B. Fund A returned a 5% profit and Fund B returned a 2% profit. How much did he invest in Fund B, if the total profit from the two funds together was $910?
Answer by ikleyn(52781)

.
Eric invested his savings in two investment funds.
The amount he invested in Fund A was 7000 less than the amount he invested in Fund B.
Fund A returned a 5% profit and Fund B returned a 2% profit.
How much did he invest in Fund B, if the total profit from the two funds together was $910?
~~~~~~~~~~~~~~

Let x be the amount invested in fund B, in dollar.

then the amount invested in fund A was (x-7000) dollars.


Write the total interest equation

    0.05(x-7000) + 0.02x = 910  dollars.


Simplify it and find x

    0.05x - 350 + 0.02x = 910

    0.05x + 0.02x = 910 + 350 

        0.07x     =    1260

            x     =    1260/0.07 = 18000.


ANSWER.  The amount invested in fund B was 18000 dollars.

Solved.